# Diagnostically Competitive Performance of a Physiology-Informed Generative Multi-Task Network for Contrast-Free CT Perfusion

### v1

@@ Abstract
Perfusion imaging is extensively utilized to assess hemodynamic status and tissue perfusion in various organs. Computed tomography perfusion (CTP) imaging plays a key role in the early assessment and planning of stroke treatment. While CTP provides essential perfusion parameters to identify abnormal blood flow in the brain. However, the use of contrast agents in CTP can lead to allergic reactions and adverse side effects, along with costing \$4.9 billion worldwide in 2022. To address these challenges, we propose a novel deep learning framework called Multitask Automated Generation of Intermodal CT perfusion maps (MAGIC). This framework combines generative artificial intelligence and physiological information to map non-contrast computed tomography (CT) imaging to multiple contrast-free CTP imaging maps. We demonstrate enhanced image fidelity by incorporating physiological characteristics into the loss terms. Our network was trained and validated using CT image data from patients referred for stroke at UF Health and demonstrated robustness to abnormalities in brain perfusion activity. A double-blinded study was conducted involving seven experienced neuroradiologists and vascular neurologists. This study validated MAGIC’s visual quality and diagnostic accuracy showing favorable performance compared to clinical perfusion imaging with intravenous contrast injection. Overall, MAGIC holds great promise in revolutionizing healthcare by offering contrast-free, cost-effective, and rapid perfusion imaging.
Keywords Generative adversarial network  $\cdot$ Non-contrast perfusion  $\cdot$ Non-contrast CT  $\cdot$ Diagnostic evaluation  $\cdot$ Ischemic stroke  $\cdot$ Brain perfusion
@@ 1 Introduction
Stroke is a second leading cause of death in the world [30] with more than 795,500 strokes occur in the U.S. each year [31]. Prompt treatment of stroke patients has been associated with significantly improved outcomes, allowing patients to return to their families and resume normal life. Therefore, identifying and characterizing the hypo-perfused tissue in acute ischemic stroke is crucial to ensuring the best clinical outcomes [32]. Non-contrast computed tomography (NCCT) is used to exclude diagnoses of brain tumors and intracerebral hemorrhages but other advanced modalities such as CT angiography and CTP are necessary to accurately identify patients with acute ischemic stroke [33]. Up to 40% of stroke patients have a normal NCCT scan within the first few hours [34] of stroke, highlighting the need of CTP in assessing cerebral perfusion and its value in identifying potentially salvageable brain tissue [35]. CTP imaging works by imaging a volume of brain tissue during the intravenous injection of an iodinated contrast bolus. This data is used to characterize the capillary-level perfusion within the brain and identify any regions of abnormal blood flow and blood volume. Thus, CTP is important for evaluating the stroke status and informing therapeutic decision making. Prompt therapy can then be used to restore blood flow to the tissue through the use of tissue plasminogen activator or thrombectomy [17, 36]. Additionally, the mismatch between cerebral blood flow / cerebral blood volume or mean transit time / cerebral blood volume can be used to identify salvageable brain tissue that may recover. However, even though CTP is beneficial, this imaging procedure is lengthy, requires contrast agent injection, and may delay stroke treatment. Moreover, many healthcare settings—especially rural or low-resource areas—do not have access to perfusion-capable CT scanners, and not all CT scanners support CTP protocols [37]. Mobile stroke units, which are becoming increasingly common, also typically rely on standard NCCT rather than full CTP due to equipment and time constraints. This lack of access can hinder timely decision-making for interventions such as thrombectomy, especially when preparing for the operating room where every minute counts. Having CT perfusion maps generated by AI methods could significantly accelerate treatment and improve outcomes. Therefore, we propose the use of generative artificial intelligence and physiological information to map NCCT imaging to contrast-free CTP imaging maps.
@@ 2.1.1 Data acquisition
Neuroimaging data were retrospectively retrieved at UF Health between 2013 and 2022 with Institutional Review Board (IRB) approval by the University of Florida. Written informed consent was waived by the IRB. The perfusion studies were conducted using an injection of an iodine-based contrast agent, and the collected data were analyzed using RAPID CTP analysis software (iSchemaView, Menlo Park, CA., USA) to generate each CTP map. The data were queried to include patients whose imaging data included RAPID CTP data. 880 patients were found to have both RAPID CTP and NCCT imaging data. 39 patients were excluded due to incomplete data present, resulting in 841 patients being selected for this study. Among them, 673 patients (80%) were randomly selected for training. The remaining 168 patients (20%) were reserved for testing, and 20 patients from this test set were randomly selected for diagnostic evaluation. Furthermore, we statically applied data augmentation to the training data [38]: random horizontal reflection, random rotation between -10 degrees and 10 degrees, and random vertical and horizontal translations between -30 and 30 pixels in each direction [39]. After extracting and applying augmentations to paired NCCT and CTP data, 29,358 (26.8 GB) paired slices of NCCT and CTP images were used to train MAGIC. Each image was normalized with a mean of 0 and a standard deviation of 1 in the training environment.
@@ 2.1.3 Ischemia severity
The patients in this dataset presented a variety of stroke presentations, ranging from healthy to severely infarcted. RAPID yields diagrams for each patient of identified regions of infarct core (defined by CBF $<$ 30%) and ischemic penumbra (defined by TTP $>$ 6s)[40]. Studies have shown that these are accurate surrogate regions for the infarct core and ischemic penumbra regions, which we use as surrogates for the infarct core and ischemic penumbra regions below. The indicated regions in the RAPID data were used to create ROI masks to further evaluate MAGIC’s performance in encoding infarcted brain regions. To evaluate the infarction severity of each subject, the ratio of infarct core and ischemic penumbra was defined as the ratio of infarct core to ischemic penumbra in the 3-dimensional volume of the whole brain. This ratio is also called RAPID Ratio or the Mismatch Ratio (MMR) [41, 42, 39, 43]. MMR was calculated by summing the number of pixels of reduced CBF and delayed TTP volumes respectively for each axial slices in the CBF/Mismatch analysis. These ratios were used to classify each patient as either presenting no infarction, or with different levels of infarction. The class of no infarction represents core-to-penumbra ratio of 0, and with infarction group has a core-to-penumbra ratio greater than 0. Among the infarction group, mild to moderate infarction is defined by a core-to-penumbra ratio larger than 0 but less than or equal to 0.5, and severe group has a ratio of greater than 0.5. These ratio classes were determined with the inspiration of Straka et al.[44] and exactly chosen with the help of the radiologists.
@@ 2.2.1 Perfusion maps
A novel GAN architecture is proposed for the task of NCCT to CTP translation. MAGIC is based on the traditional pix2pix architecture with novel multitask architecture and physiology-informed loss functions as an image-to-image translation network [47]. In contrast to traditional image translation networks, MAGIC enables greater performance in this task due to 1) simultaneous generation of multiple CTP maps, 2) shared encoding layers between perfusion maps to improve performance and reduce training costs, 3) physiology-informed loss terms that leverage MAGIC’s multitask architecture, 4) real-time generation of perfusion maps without contrast agent injection, 5) high spatial resolution in the axial direction and 3D models, 6) generalizability to any additional perfusion maps, and 7) a novel Physician-In-The-Loop (PILO) module, which is a tunable convolution layer to provide a physician-computer interface to allow physicians adjust the contribution of perfusion and anatomical information in the synthesized CTP maps to facilitate clinical decision making. The overview of the proposed method is presented in Fig. 1.
Figure 1: Proposed MAGIC pipeline for non-contrast CT perfusion map generation. a. Neighboring NCCT slices in the z-axis are stacked to create a pseudo-RGB image and are matched with the corresponding perfusion maps to construct paired datasets. b. MAGIC consists of a generative adversarial network to co-learn between the generator and discriminator networks using novel physiology-informed loss terms. During training, the generator synthesizes perfusion maps using a UNet architecture, which are refined through adversarial learning with the PatchGAN discriminator. In the testing stage — indicated by the dashed line in the figure — the Physician-in-the-Loop (PILO) interactive module allows clinicians to visually assess and adjust the optimal ratio between non-contrast structural information and synthesized perfusion information. c. The composite loss terms contribute synergistically to MAGIC’s high-quality performance. During training, GAN loss (LGAN) and L1 loss (LL1) are used to ensure style consistency and structural fidelity in the synthesized imaging. The novel extrema loss (Lext) is used to ensure proper encoding of ischemic tissue regions. The novel multimodal loss (LMML) is used to ensure physiological consistency across the synthesized perfusion maps. d. It was determined that seven experienced neuroradiologists could not differentiate the synthetic and real CT perfusion imaging in a double-blinded review in terms of the realistic level and diagnostic quality.
@@ 2.3.1 Notations
We adopt a similar notation to Zhu et al [6]. The goal of our modal is to learn four mapping functions between the domain X of NCCT images and the domain of perfusion maps (CBV, CBF, MTT, and TTP). Given M training samples, each is composed of five elements: ${\{x_{i},cbv_{i},cbf_{i},mtt_{i},ttp_{i}\}}_{i}^{M}$ for the $i^{th}$ sample. Here $x_{i}\in X$ where X represents the set of NCCT images, $cbv_{i}\in CBV$ , $cbf_{i}\in CBF$, $mtt_{l}\in MTT$, and $ttp_{m}\in TTP$. We will refer to the set containing all four perfusion maps as $CTP={CBV,CBF,MTT,TTP}$. Let N denote the cardinality of CTP which is 4 in our case but can be flexible depending on how many perfusion maps to predict for the generalizability of this method. Recalling that X contains the real NCCT, we denote the input data distribution as $x~{}p_{data}(x)$ and will indicate the data distribution of the different perfusion maps as $cbv\sim p_{data}(cbv)$,$cbf\sim p_{data}(cbf)$,$mtt\sim p_{data}(mtt)$, and $ttp\sim p_{data}(ttp)$. Importantly, we denote the joint data distribution of an NCCT and its corresponding real perfusion map as $(x,map)\sim(p_{data}(x),p_{data}(map))$ where map can be replaced by any type from ${cbv,cbf,mtt,ttp}$ depending on the perfusion map being discussed. The model includes four mappings represented by four generators: $G_{CBV}:X\rightarrow CBV$, $G_{CBF}:X\rightarrow CBF$, $G_{MTT}:X\rightarrow MTT$, and $G_{TTP}:X\rightarrow TTP$. Moreover, the model has four discriminators $D_{CBV}$,$D_{CBF}$,$D_{MTT}$,and $D_{TTP}$ where each discriminator aims to distinguish between its corresponding real perfusion map and the synthetic perfusion map. Finally, we use $\odot$ to represent element-wise multiplication and $W_{x}$ to represent a weighted matrix of a perfusion map of a given input x (real NCCT image). The weight matrices are constructed by taking a real perfusion map image, normalizing the pixel values, and squaring so all pixel values are positive.
@@ 2.3.2 Physiology-Informed Loss Functions
$\begin{array}[]{c}\mathcal{L}_{\text{GAN}}(G_{CTP},D_{CTP},CTP,X)=\\ \frac{1}{N}\sum_{PM\in CTP}\mathbb{E}_{PM\sim p_{\text{data}}(PM)}\left[\log D_{PM}(PM)\right]\\ +\mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\log(1-D_{PM}(G_{PM}(x)))\right]\end{array}$ (1)
$\begin{array}[]{c}\mathcal{L}_{L1}(G_{CTP},CTP,X)=\\ \frac{1}{N}\sum_{PM\in CTP}\mathbb{E}_{(x,PM)\sim(p_{\text{data}}(x),p_{\text{data}}(PM))}\left[\|PM-G_{PM}(x)\|_{1}\right]\end{array}$ (2)
We propose two novel loss terms, multimodal loss and extrema loss, in MAGIC to leverage the physiological rules between CTP map types to further improve performance. Specifically, the multimodal loss leverages the Central Volume Principle [49], which describes the physiological relationship between the CBF, CBV, and MTT map classes. These map types are related by Eq. (3), which leads to the multimodal loss represented in Eq. (4). The generator minimizes the L1 distance between the real CBV map and the product of the synthesized CBF and MTT maps, ensuring that the synthetic CTP maps are consistent with the central volume principle.
@@
$\begin{array}[]{c}\mathcal{L}_{\text{MML}}(G_{CBF},G_{MTT},CBV,X)=\\ \mathbb{E}_{(x,CBV)\sim(p_{\text{data}}(x),p_{\text{data}}(CBV))}\left[\|G_{CBF}(x)\times G_{MTT}(x)-CBV\|_{1}\right]\end{array}$ (4)
The extrema loss term enables increased fidelity in encoding regions-of-interest (ROI) in the simulated scans that differ significantly from the mean of the observed image. The extrema loss was introduced to ensure that MAGIC would accurately generate the ROIs with both very high and very low pixel intensities compared to the slice mean value. Its inclusion increases the likelihood that stroke regions in the synthesized perfusion imaging are properly identified and learned by MAGIC during training. Infarcted tissue regions tend to exhibit elevated MTT and TTP and decreased CBF and CBV, so these characteristics must be present in the synthetic CTP maps to ensure acceptable diagnostic quality. We generated a weight map $W_{x}$ by normalizing the real perfusion map between [-0.5, 0.5] and then take an element-wise square of the normalized image, as shown in Eq. (5). $W_{x}$ emphasizes the ROIs with values farther from the image mean. Then this weight map is multiplied by the MSE between the real and synthesized perfusion imaging, shown in Eq. (6). The mean of this product is taken to produce the final extrema loss term, as provided in Eq. (7). The weight matrices are constructed by taking a real perfusion map image, normalizing the pixel values, and squaring so all pixel values are positive.
$W_{x}=(\frac{PM-min(PM)}{max(PM)-min(PM)}-0.5)^{2}$ (5)
$H_{x}=(\frac{G_{PM}(x)-min(G_{PM}(x))}{max(G_{PM}(x))-min(G_{PM}(x))}-\frac{PM-min(PM)}{max(PM)-min(PM)})^{2}$ (6)
@@
We weigh each additional loss term with lambda tuning hyperparameters. The final objective we aim to solve, which $G_{CTP}$ aims to minimize and $D_{CTP}$ aims to maximize, is shown below in Eq. (8).
$\begin{array}[]{c}{{G^{\ast}=arg\min\limits_{G_{CTP}}\max\limits_{D_{CTP}}\mathcal{L}_{GAN}(G_{CTP},D_{CTP},CTP,X)}}\\ {{+\lambda_{1}\mathcal{L}_{L1}(G_{CTP},CTP,X)}}\\ {{+\lambda_{2}\mathcal{L}_{EXT}(G_{CTP},CTP,X)}}\\ {{+\lambda_{3}\mathcal{L}_{MML}(G_{CBF},G_{MTT},CBV,X)}}\end{array}$ (8)

### v2

@@ Abstract
Perfusion imaging is extensively utilized to assess hemodynamic status and tissue perfusion in various organs. Computed tomography perfusion (CTP) imaging plays a key role in the early assessment and planning of stroke treatment. While CTP provides essential perfusion parameters to identify abnormal blood flow in the brain. However, CTP can be expensive with limited accessibility, and the use of contrast agents in CTP can lead to allergic reactions and adverse side effects. To address these challenges, we propose a novel deep learning framework called Multitask Automated Generation of Intermodal CT perfusion maps (MAGIC). This framework combines generative artificial intelligence and physiological information to map non-contrast computed tomography (CT) imaging to multiple contrast-free CTP imaging maps. We demonstrate enhanced image fidelity by incorporating physiological characteristics into the loss terms. Our network was trained and validated using CT image data from patients referred for stroke at UF Health and demonstrated robustness to abnormalities in brain perfusion activity. A double-blinded study was conducted involving seven experienced neuroradiologists and vascular neurologists. This study validated MAGIC’s visual quality and diagnostic accuracy showing favorable performance compared to clinical perfusion imaging with intravenous contrast injection. Overall, MAGIC holds great promise in revolutionizing healthcare by offering contrast-free, cost-effective, and rapid perfusion imaging.
keywords:
, Generative adversarial network , Non-contrast perfusion , Non-contrast CT , Diagnostic evaluation , Ischemic stroke , Brain perfusion
[…]
@@ 1 Introduction
Stroke is a second leading cause of death in the world [38] with more than 795,500 strokes occur in the U.S. each year [37]. Prompt treatment of stroke patients has been associated with meaningful improved outcomes, allowing patients to return to their families and resume normal life. Therefore, identifying and characterizing the hypo-perfused tissue in acute ischemic stroke is crucial to ensuring the best clinical outcomes [50]. Non-contrast computed tomography (NCCT) is used to exclude diagnoses of brain tumors and intracerebral hemorrhages but other advanced modalities such as CT angiography and CTP are necessary to accurately identify patients with acute ischemic stroke [59]. Up to 40% of stroke patients have a normal NCCT scan within the first few hours [36] of stroke, highlighting the need of CTP in assessing cerebral perfusion and its value in identifying potentially salvageable brain tissue [5]. CTP imaging works by imaging a volume of brain tissue during the intravenous injection of an iodinated contrast bolus. These data are used to characterize the capillary-level perfusion within the brain and identify any regions of abnormal blood flow and blood volume. Thus, CTP is important for evaluating the stroke status and informing therapeutic decision making. Prompt therapy can then be used to restore blood flow to the tissue through the use of tissue plasminogen activator or thrombectomy [24, 14]. Additionally, the mismatch between cerebral blood flow / cerebral blood volume or mean transit time / cerebral blood volume can be used to identify salvageable brain tissue that may recover. However, even though CTP is beneficial, this imaging procedure is lengthy, requires contrast agent injection, and may delay stroke treatment. Moreover, many healthcare settings—especially rural or low-resource areas—do not have access to perfusion-capable CT scanners, and not all CT scanners support CTP protocols [4]. Mobile stroke units, which are becoming increasingly common, also typically rely on standard NCCT rather than full CTP due to equipment and time constraints. This lack of access can hinder timely decision-making for interventions such as thrombectomy, especially when preparing for the operating room where every minute counts. Having CT perfusion maps generated by AI methods could assist clinicians to accelerate treatment and improve outcomes. Therefore, we propose the use of generative artificial intelligence and physiological information to map NCCT imaging to contrast-free CTP imaging maps.
@@ 2.1.1 Data acquisition
Neuroimaging data were retrospectively retrieved at UF Health between 2013 and 2022 with Institutional Review Board (IRB) approval by the University of Florida. Written informed consent was waived by the IRB. The perfusion studies were conducted using an injection of an iodine-based contrast agent, and the collected data were analyzed using RAPID CTP analysis software (iSchemaView, Menlo Park, CA., USA) to generate each CTP map. The data were queried to include patients whose imaging data included RAPID CTP data. A total of 880 patients were found to have both RAPID CTP and NCCT imaging data. After excluding 39 patients due to incomplete data, 841 patients remained for analysis.
Among them, 673 patients (80%) were randomly selected for training. The remaining 168 patients (20%) were reserved as an independent test set; and 20 patients from the test set were further randomly selected for diagnostic evaluation. Furthermore, static data augmentation was applied to the training data [16]: random horizontal reflection, random rotation between -10 degrees and 10 degrees, and random vertical and horizontal translations between -30 and 30 pixels in each direction [39]. After extracting and applying augmentations to paired NCCT and CTP data, 29,358 (26.8 GB) paired slices of NCCT and CTP images were used to train MAGIC. Each image was normalized with a mean of 0 and a standard deviation of 1 in the training environment.
@@ 2.1.3 Ischemia severity
The patients in this dataset presented a variety of stroke presentations, ranging from healthy to severely infarcted. RAPID yields diagrams for each patient of identified regions of infarct core (defined by CBF << 30%) and ischemic penumbra (defined by Tmax >> 6s)[44]. Studies have shown that these are accurate surrogate regions for the infarct core and ischemic penumbra regions, which we use as surrogates for the infarct core and ischemic penumbra regions below. The indicated regions in the RAPID data were used to create ROI masks to further evaluate MAGIC’s performance in encoding infarcted brain regions. To evaluate the infarction severity of each subject, the ratio of infarct core and ischemic penumbra was defined as the ratio of infarct core to ischemic penumbra in the 3-dimensional volume of the whole brain. This ratio is also called RAPID Ratio or the Mismatch Ratio (MMR) [8, 61, 39, 60]. MMR was calculated by summing the number of pixels of reduced CBF and delayed TTP volumes respectively for each axial slices in the CBF/Mismatch analysis. These ratios were used to classify each patient as either presenting no infarction, or with different levels of infarction. The class of no infarction represents core-to-penumbra ratio of 0, and with infarction group has a core-to-penumbra ratio greater than 0. Among the infarction group, mild to moderate infarction is defined by a core-to-penumbra ratio larger than 0 but less than or equal to 0.5, and severe group has a ratio of greater than 0.5. These ratio classes were determined with the inspiration of Straka et al.[56] and exactly chosen with the help of the radiologists.
@@ 2.2.1 Perfusion maps
A novel GAN architecture is proposed for the task of NCCT to CTP translation. MAGIC is based on the traditional pix2pix architecture with novel multitask architecture and physiology-informed loss functions as an image-to-image translation network [19]. In contrast to traditional image translation networks, MAGIC enables greater performance in this task due to 1) simultaneous generation of multiple CTP maps, 2) shared encoding layers across all perfusion maps, allowing the generator to extract common anatomical features and reduce the number of trainable parameters, 3) physiology-informed loss terms that leverage MAGIC’s multitask architecture, 4) real-time generation of perfusion maps without contrast agent injection, 5) high spatial resolution in the axial direction and 3D models, 6) generalizability to any additional perfusion maps, and 7) a novel Physician-In-The-Loop (PILO) module, which is a tunable convolution layer to provide a physician-computer interface to allow physicians adjust the contribution of perfusion and anatomical information in the synthesized CTP maps to facilitate clinical decision making. The overview of the proposed method is presented in Fig. 1.
Figure 1: Proposed MAGIC pipeline for non-contrast CT perfusion map generation. a. Neighboring NCCT slices in the z-axis are stacked to create a pseudo-RGB image and are matched with the corresponding perfusion maps to construct paired datasets. b. MAGIC consists of a generative adversarial network to co-learn between the generator and discriminator networks using novel physiology-informed loss terms. During training, the generator synthesizes perfusion maps using a UNet architecture, which are refined through adversarial learning with the PatchGAN discriminator. In the testing stage — indicated by the dashed line in the figure —the Physician-in-the-Loop (PILO) interactive module allows clinicians to visually assess and adjust the optimal ratio between non-contrast structural information and synthesized perfusion information.
c. The composite loss terms contribute synergistically to MAGIC’s high-quality performance. During training, GAN loss (LGAN) and L1 loss (LL1) are used to ensure style consistency and structural fidelity in the synthesized imaging. The novel extrema loss (Lext) is used to ensure proper encoding of ischemic tissue regions. The novel multimodal loss (LMML) is used to ensure physiological consistency across the synthesized perfusion maps. d. It was determined that seven experienced neuroradiologists could not differentiate the synthetic and real CT perfusion imaging in a double-blinded review in terms of the realistic level and diagnostic quality.
@@ 2.3.1 Notations
We adopt a similar notation to Zhu et al [69]. The goal of our modal is to learn four mapping functions between the domain X of NCCT images and the domain of perfusion maps (CBV, CBF, MTT, and TTP). Given M training samples, each is composed of five elements: ${\{x_{i},cbv_{i},cbf_{i},mtt_{i},ttp_{i}\}}_{i}^{M}$ for the $i^{th}$ sample. Here $x_{i}\in X$ where X represents the set of NCCT images, $cbv_{i}\in CBV$ , $cbf_{i}\in CBF$, $mtt_{l}\in MTT$, and $ttp_{m}\in TTP$. We will refer to the set containing all four perfusion maps as CTP=CBV,CBF,MTT,TTPCTP={CBV,CBF,MTT,TTP}. Let N denote the cardinality of CTP which is 4 in our case but can be flexible depending on how many perfusion maps to predict for the generalizability of this method. Recalling that X contains the real NCCT, we denote the input data distribution as $x~p_{data}(x)$ and will indicate the data distribution of the different perfusion maps as $cbv\sim p_{data}(cbv)$$cbf\sim p_{data}(cbf)$$mtt\sim p_{data}(mtt)$, and $ttp\sim p_{data}(ttp)$. Importantly, we denote the joint data distribution of an NCCT and its corresponding real perfusion map as $(x,map)\sim(p_{data}(x),p_{data}(map))$ where map can be replaced by any type from cbv,cbf,mtt,ttp{cbv,cbf,mtt,ttp} depending on the perfusion map being discussed. The model includes four mappings represented by four generators: $G_{CBV}:X\rightarrow CBV$, $G_{CBF}:X\rightarrow CBF$, $G_{MTT}:X\rightarrow MTT$, and $G_{TTP}:X\rightarrow TTP$. Moreover, the model has four discriminators $D_{CBV}$$D_{CBF}$$D_{MTT}$,and $D_{TTP}$ where each discriminator aims to distinguish between its corresponding real perfusion map and the synthetic perfusion map. Finally, we use $\odot$ to represent element-wise multiplication and $W_{x}$ to represent a weighted matrix of a perfusion map of a given input x (real NCCT image). The weight matrices are constructed by taking a real perfusion map image, normalizing the pixel values, and squaring so all pixel values are positive.
@@ 2.3.2 Physiology-Informed Loss Functions
$\begin{split}\mathcal{L}_{\text{GAN}}(G_{CTP},D_{CTP},CTP,X)=&\frac{1}{N}\sum_{PM\in CTP}\mathbb{E}_{PM\sim p_{\text{data}}(PM)}\left[\log D_{PM}(PM)\right]\\ &+\mathbb{E}_{x\sim p_{\text{data}}(x)}\left[\log(1-D_{PM}(G_{PM}(x)))\right]\end{split}$ (1)
$\begin{split}\mathcal{L}_{L1}(G_{CTP},CTP,X)=&\frac{1}{N}\sum_{PM\in CTP}\mathbb{E}_{(x,PM)\sim(p_{\text{data}}(x),p_{\text{data}}(PM))}\left[\|PM-G_{PM}(x)\|_{1}\right]\end{split}$ (2)
We propose two novel loss terms, multimodal loss and extrema loss, in MAGIC to leverage the physiological rules between CTP map types to further improve performance. Specifically, the multimodal loss leverages the Central Volume Principle [17], which describes the physiological relationship between the CBF, CBV, and MTT map classes. These map types are related by Eq. (3), which leads to the multimodal loss represented in Eq. (LABEL:eq:MML). The generator minimizes the L1 distance between the real CBV map and the product of the synthesized CBF and MTT maps, ensuring that the synthetic CTP maps are consistent with the central volume principle.
@@
$\begin{split}\mathcal{L}_{\text{MML}}(G_{CBF},G_{MTT},CBV,X)=&\mathbb{E}_{(x,CBV)\sim(p_{\text{data}}(x),p_{\text{data}}(CBV))}\\ &\quad\left[\|G_{CBF}(x)\times G_{MTT}(x)-CBV\|_{1}\right]\end{split}$ (4)
The extrema loss term enables increased fidelity in encoding regions-of-interest (ROI) in the simulated scans that deviate from the mean of the observed image. The extrema loss was introduced to ensure that MAGIC would accurately generate the ROIs with both very high and very low pixel intensities compared to the slice mean value. Its inclusion increases the likelihood that stroke regions in the synthesized perfusion imaging are properly identified and learned by MAGIC during training. Infarcted tissue regions tend to exhibit elevated MTT and TTP and decreased CBF and CBV, so these characteristics must be present in the synthetic CTP maps to ensure acceptable diagnostic quality. We generated a weight map $W_{x}$ by normalizing the real perfusion map between [-0.5, 0.5] and then take an element-wise square of the normalized image, as shown in Eq. (LABEL:eq:wx). $W_{x}$ emphasizes the ROIs with values farther from the image mean. Then this weight map is multiplied by the MSE between the real and synthesized perfusion imaging, shown in Eq. (6). The mean of this product is taken to produce the final extrema loss term, as provided in Eq. (7). The weight matrices are constructed by taking a real perfusion map image, normalizing the pixel values, and squaring so all pixel values are positive.
$W_{x}=\left(\frac{PM-\min(PM)}{\max(PM)-\min(PM)}-0.5\right)^{2}$ (5)
$H_{x}=\left(\frac{G_{PM}(x)-\min(G_{PM}(x))}{\max(G_{PM}(x))-\min(G_{PM}(x))}-\frac{PM-\min(PM)}{\max(PM)-\min(PM)}\right)^{2}$ (6)
@@
We weight each additional loss term with lambda tuning hyperparameters. The final objective we aim to solve, which $G_{CTP}$ aims to minimize and $D_{CTP}$ aims to maximize, is shown below in Eq. (8).
$G^{\ast}=\arg\min_{G_{CTP}}\max_{D_{CTP}}\mathcal{L}_{GAN}(G_{CTP},D_{CTP},CTP,X)$ (8)
$\quad+\lambda_{1}\mathcal{L}_{L1}(G_{CTP},CTP,X)$
$\quad+\lambda_{2}\mathcal{L}_{EXT}(G_{CTP},CTP,X)$
$\quad+\lambda_{3}\mathcal{L}_{MML}(G_{CBF},G_{MTT},CBV,X)$